# LLM-Based Program Generation for Triggering Numerical Inconsistencies Across Compilers

### v1

@@ 3.2.3. Evaluation Results
Table 2 compares the effectiveness of LLM4FP and the baselines at generating floating-point programs that trigger numerical inconsistencies in terms of the metrics described in Section 3.2.2. Among all approaches, LLM4FP achieves the highest inconsistency rate at 26.56%, more than doubling that of Varity, 11.93%. In total, LLM4FP discovers 4,781 inconsistencies among the total 18,000 comparisons across compilers and optimization levels. This indicates that incorporating the program structural guidance via the Varity grammar specification and a feedback loop via examples of inconsistent programs enables the LLM to generate test programs that are more likely to trigger numerical inconsistencies across compilers and/or optimization levels. Grammar-Guided also outperforms Varity and Direct-Prompt by 3.87% and 2.37%, respectively. This suggests that grammar-guided generation combined with the LLM alone improves inconsistency detection.
@@
In terms of the time cost, Varity remains the most efficient approach requiring only about 30 minutes to complete the full pipeline for 1,000 generated programs, while all LLM-based approaches take approximately 3 hours. The additional cost arises from the latency of invoking the LLM API during program generation, which accounts for around 30% of the total time. In the future, this overhead could potentially be reduced by batching API calls or leveraging local LLM inference to lower latency. Though LLM4FP has a longer runtime than Varity, its improvement in inconsistency detection indicates a trade-off between testing effectiveness and time cost.
Table 3. Comparing LLM4FP with baselines in terms of program diversity. Lower CodeBLEU score is better. Clone Types 1, 2, 2c: Type-1, Type-2, and Type-2c code clones.
Approach CodeBLEU Clone Types Percentage
1 2 2c
Varity 0.3581 0 0 0 0.00%
Direct-Prompt 0.4213 0 0 0 0.00%
Grammar-Guided 0.5099 0 14 6 2.00%
LLM4FP 0.3610 4 36 22 6.20%
Because it is important to avoid repeatedly triggering the same numerical inconsistencies, diversity in program generation is crucial. Table 3 compares LLM4FP with the baselines in terms of the CodeBLEU score among the generated programs. A lower score is better, as it indicates more diverse programs. The results show that Varity achieves the best diversity with a score of 0.3581, though LLM4FP follows closely with a score of 0.3610 that is only 0.0029 larger. Compared to Varity, Direct-Prompt increases the CodeBLEU score by 17.65%, reflecting reduced diversity. We infer that when given open-ended prompts, the LLM tends to follow safe and common patterns to generate programs. Grammar-Guided increases the CodeBLEU score substantially by 42.39% over Varity. When provided with a grammar specification, the LLM tends to adhere to the given structure and reuse similar syntactic patterns across generations. This shows a limitation of LLMs: strict constraints make them repeat patterns instead of exploring diverse solutions. By contrast, LLM4FP clearly improves program diversity over the other two LLM variants. By showing examples of previously successful programs and explicitly instructing the LLM to generate different ones, the model is guided to avoid reusing previous structures while still learning which patterns are effective at triggering numerical inconsistencies.
Table 3 also reports the number of Type-1, Type-2, and Type-2c clones, along with the percentage of generated programs that are clones, which may indicate redundancy and little additional value. LLM4FP generates the most number of code clones across the three clone types, while Varity and Direct-Prompt produce none. However, the number of clones detected in LLM4FP is still relatively small, accounting for 6.20% of all programs. We consider this acceptable given the total number of inconsistencies our approach finds, shown in  Table 2.
@@
Response to RQ1: LLM4FP achieves the highest inconsistency rate at 26.56%. This comes to a higher cost due to the latency of API LLM calls, and to the presence of a small number of potentially redundant programs. On the other hand, Grammar-Guided and Direct-Prompt’s inconsistency rates are 15.8% and 13.43%, while having similar time cost as LLM4FP. Finally, Varity achieves the lowest inconsistency rate of 11.93% but also remains the most efficient approach both in terms of time cost and program diversity.
3.3. RQ2: Characterizing Numerical Inconsistencies
@@ 3.3.1. Experimental Setup
Since LLM4FP outperforms the other LLM-based variants, we focus the rest of the evaluation on comparing LLM4FP with Varity with respect to the numerical inconsistencies triggered by their generated floating-point programs. In this RQ, we discuss how often different classes of numerical inconsistencies are triggered. We follow Varity to investigate five classes of numerical inconsistencies: (normal and subnormal numbers), Zero (positive zero and negative zero), +Inf (positive infinity), -Inf (negative infinity), and NaN (not-a-number). Specifically, we count the number of times that there is a numerical inconsistency of the kind $\{\textit{$r_{i}$},\textit{$r_{j}$}\}$, where $_{i}\neq r$j and both results belong to different numerical categories in $\{,\textit{Zero},\textit{+Inf},\textit{-Inf},\textit{NaN}\}$. For example, an inconsistency between a real number and a zero is counted once as a {, Zero } inconsistency.
@@ 3.3.2. Evaluation Results
Optimization Level , , Zero , +Inf , -Inf
O0_nofma 605 52
O0 705 52 2
O1 759 52
O2 759 52
O3 610 52
O3_fastmath 977 100 2 2
Total Inconsistencies 4,781
@@
When further breaking down each inconsistency kind by optimization level for LLM4FP, we find that the $\{,\}$ and the $\{,\textit{Zero}\}$ inconsistency classes appear across all optimization levels, as shown in Table 4. For both kinds, O3_fastmath contributes the largest number of inconsistencies. The extreme values from $\{,\textit{+Inf}\}$ and $\{,\textit{-Inf}\}$ are extremely rare and only occur under O3_fastmath and occasionally at O0. The results show that the number of inconsistencies across optimization levels is relatively stable, except for O3_fastmath, which consistently amplifies inconsistencies.
Response to RQ2: The most common numerical inconsistency triggered by the programs generated by LLM4FP is the $\{,\}$ kind, which accounts for over 92% of all inconsistencies and appears across all optimizations. In contrast, extreme-value categories involving +Inf, -Inf, or NaN are rare and occur in the most aggressive optimization level, O3_fastmath.
@@ 3.4. RQ3: Inconsistencies Across Compilers
In this RQ, we analyze whether numerical inconsistencies are more likely to occur between different host compilers or between host and device compilers, and how optimization levels influence these inconsistencies. We also analyze all possible combinations of compilers to see which combinations produce the largest and smallest differences. We consider the 16 first floating-point digits, and report an inconsistency if any of them do not match. Table 5 summarizes the inconsistency rates for LLM4FP and Varity across compiler pairs and optimization levels. In addition to the rates, it also reports the minimum, maximum, and average number of differing digits observed in inconsistent results.
@@
For both approaches, the highest inconsistency rates consistently happen in comparisons between host and device compilers, i.e., gcc vs. nvcc and clang vs. nvcc, rather than between host compilers such as gcc vs. clang. For example, in LLM4FP, the most inconsistent pair is gcc vs. nvcc, with a total inconsistency rate of 11.71%, compared to 8.00% for gcc vs. clang. A similar trend is observed in Varity, where the highest rate of 5.50% is from clang vs. nvcc, much higher than that of gcc vs. clang which is 1.17%.
@@
Optimization level plays an important role in how frequent and when numerical inconsistencies occur. While similar trends were observed in  Section 3.3 across optimization levels, here we focus specifically on inconsistencies between host and device compilers. For Varity, the highest inconsistency rates consistently occur at the O3_fastmath level, reaching up to 2.06% for clang vs. nvcc, while the rates at other optimization levels remain very low, mostly below 1%. This is expected as aggressive optimizations and fast-math transformations can increase the chance of observing inconsistencies across compilers. For LLM4FP, O3_fastmath is also the highest for most cases, e.g., gcc vs. nvcc at 2.67%, but other optimization levels show relatively substantial inconsistency rates as well. For example, for clang vs. nvcc, the highest inconsistency rate, 2.07%, is at O0. This suggests that LLM4FP triggers inconsistencies more broadly across different optimization levels, not just under aggressive optimizations. From this, we can see the potential of LLM4FP to generate more complex and varied floating-point programs.
Beyond inconsistency rates, the digit difference results provide additional insight into the severity of numerical discrepancies. For both Varity and LLM4FP, the minimum number of differing digits is typically very small, e.g., often just 1, showing that in many cases inconsistencies are minor perturbations. However, the maximum differences can span the full 16-digit precision of double-precision values, especially in LLM4FP where all three compiler pairs reach the maximum of 16 digits at O0_nofma, O0, and O3_fastmath. On average, LLM4FP inconsistencies involve larger digit differences which are mostly above 10 digits compared to Varity, where averages are typically below 5 digits except under O3_fastmath. This indicates that LLM4FP not only shows higher rates of inconsistencies, but those inconsistencies are also much larger in terms of numerical error.
Response to RQ3: Numerical inconsistencies are observed for both host-device and host-host compiler comparisons, but are more frequent between host-device compilers. The most aggressive optimization, O3_fastmath, causes the largest number of inconsistencies. Compared to Varity, LLM4FP discovers more inconsistencies even at lower optimization levels, indicating its ability to expose numerical differences beyond those revealed by aggressive optimizations.
@@ 3.5. RQ4: Comparing to O0 without FMA
Table 6. Inconsistency rates between any optimization level and O0_nofma for Varity and LLM4FP (higher inconsistency rates are preferred)
@@
As shown in Table 6, the inconsistency rate varies when comparing O0_nofma with different optimization levels. For Varity, differences at lower optimization levels from O1 to O3 are almost zero, i.e., at most 0.04%. It mainly detects differences caused by the most aggressive optimization level, O3_fastmath, especially for clang which shows an inconsistency rate of 2.07% compared to O0_nofma. However for LLM4FP, there are also noticeable inconsistencies from O1 to O3, e.g., 1.95% for clang. The results indicate that LLM4FP’s programs are better at finding variations even when optimizations are less aggressive. For both Varity and LLM4FP, clang has the highest number of numerical inconsistencies across its optimization levels, while nvcc is the most stable. Compared to Varity, LLM4FP finds more variations across optimization levels within each compiler.
Response to RQ4: Different optimizations within the same compiler can introduce different numerical inconsistencies compared to -O0_nofma. LLM4FP finds inconsistencies across all optimization levels, not just the most aggressive ones. clang has the most inconsistencies while nvcc is the most stable.
@@ 4. Limitations
While LLM4FP improves the state-of-the-art in generating floating-point programs that trigger numerical inconsistencies, it also has several limitations. First, as indicated by CodeBLEU scores, although the feedback mechanism of LLM4FP improves the program diversity over other LLM-based variants, our approach does not show an improvement compared to Varity. We believe this is partly because we focus more on syntax guidance rather than on deeper semantic exploration. Future work will focus on incorporating semantic guidance to improve diversity and discover more numerical inconsistencies. Second, our approach has a higher runtime cost due to the latency of LLM API calls which may limit scalability for large-scale testing. Nevertheless, the improved effectiveness of LLM4FP helps offset this overhead. Third, while we explicitly instruct the LLM to avoid undefined behavior, we cannot guarantee its absence, which may cause false positives in inconsistency detection. In the future, we plan to integrate undefined behavior checkers or static analysis techniques to prevent invalid programs. Fourth, the effectiveness of LLM4FP depends heavily on carefully crafted prompts. If the prompts are poorly designed, the effectiveness of triggering numerical inconsistencies will be limited. Lastly, not all inconsistencies detected by LLM4FP indicate real bugs. Some may reflect allowed differences in compiler behavior, while others can impact program stability and lead to incorrect results. Future work should focus on distinguishing harmful bugs from acceptable floating-point inconsistencies.
@@ 5. Related Work
Compiler testing has been widely studied to ensure the correctness and reliability of compiler behavior (Chen et al., 2021a). Traditional compiler testing techniques include random testing (Duran and Ntafos, 1984; Ince and Hekmatpour, 1986), randomized differential testing (McKeeman, 1998; Groce et al., 2007), equivalence modulo inputs testing (Le et al., 2014), program enumeration (Zhang et al., 2017) among others. The main challenge in compiler testing is to generate diverse, syntactically and semantically valid programs to test the compiler behaviors. To address it, researchers have proposed various test program generation approaches. Grammar-based approaches synthesize programs from scratch by following generation rules to ensure syntactic validity, such as CSmith (Yang et al., 2011), YARPGen (Livinskii et al., 2020), and LangFuzz (Holler et al., 2012). To generate more realistic and complicated test programs, mutation-based approaches (Donaldson et al., 2017; Le et al., 2014, 2015; Zhang et al., 2017) are proposed to mutate valid seed programs for generating new ones.
@@
Varity (Laguna, 2020) extends compiler testing to the domain of floating-point computation, focusing not on compiler bugs but on numerical inconsistencies across different compilers and architectures. It uses grammar-guided random generation to synthesize test programs but does not leverage domain knowledge or feedback from previous tests, which limits its ability to discover floating-point inconsistencies efficiently. In contrast, our work uses LLMs to generate and adapt test programs based on prior results, enabling us to discover more numerical inconsistencies across compilers.
@@
Recent advances in LLMs have opened new opportunities for test program generation (Jiang et al., 2024). LLMs like Codex (Chen et al., 2021b) and GPT (OpenAI, 2023) have demonstrated strong capabilities in generating syntactically correct and semantically meaningful code from natural language prompts in tools like TestPilot (Schäfer et al., 2023) and ChatTester (Yuan et al., 2023). Several recent works have applied LLMs to testing compilers (Xia et al., 2024; Yang et al., 2024; Munley et al., 2024; Gu, 2023). For example, WhiteFox (Yang et al., 2024) integrates LLMs with compiler source code to guide white-box fuzzing for optimization testing. Fuzz4All (Xia et al., 2024) uses LLMs to fuzz compilers across multiple languages via auto-prompting and iterative mutation. Unlike our work, they focus on general bug detection and do not target floating-point inconsistencies. Overall, our work bridges the gap between LLM-based program generation and numerical inconsistency detection across compilers, which is underexplored in prior research.

### v2

@@ 3.2.3. Evaluation Results
[…]
@@
In terms of the time cost, Varity remains the most efficient approach requiring only about 30 minutes to complete the full pipeline for 1,000 generated programs, while LLM-based approaches take approximately from 4 to 6 hours. The additional cost arises from the latency of invoking the LLM API during program generation, which accounts for more than half of the total time. In the future, this overhead could potentially be reduced by batching API calls or leveraging local LLM inference to lower latency. Though LLM4FP has a longer runtime than Varity, its improvement in inconsistency detection indicates a trade-off between testing effectiveness and time cost.
Because it is important to avoid repeatedly triggering the same numerical inconsistencies, diversity in program generation is crucial. Table 2 compares LLM4FP with the baselines in terms of the CodeBLEU score among the generated programs. A lower score is better, as it indicates more diverse programs. The results show a clear descending trend in CodeBLEU scores from Varity (0.3581) to Direct-Prompt (0.3561), Grammar-Guided (0.3442), and finally to LLM4FP (0.2788), suggesting that program diversity consistently improves across these approaches. Among them, LLM4FP achieves the highest diversity, reducing the CodeBLEU score by 22.14% compared to Varity. We attribute this improvement to the feedback-driven generation mechanism of LLM4FP, which iteratively mutates one existing example based on the model’s prior successes. This feedback loop encourages the LLM to explore new program variants that differ behaviorally from previous ones while retaining effective structures for triggering numerical inconsistencies. By explicitly guiding the model to refine and diversify rather than simply regenerate, LLM4FP improves the diversity of program generation while maintaining meaningful program structures.
Regarding Type-1, Type-2, and Type-2c clones, none were detected for any approach, showing that the generated programs across all methods are structurally distinct. This confirms that the prompt design in our LLM-based approaches effectively avoids trivial code duplication while maximizing the discovery of unique inconsistencies.
[…]
@@
Response to RQ1: LLM4FP achieves the highest inconsistency rate at 29.33% and the best program diversity with the lowest CodeBLEU score of 0.2788. This comes to a higher cost due to the latency of API LLM calls. On the other hand, Grammar-Guided and Direct-Prompt’s inconsistency rates are 16.47% and 14.23%, while having similar time cost as LLM4FP. Finally, Varity achieves the lowest inconsistency rate of 11.93% but also remains the most efficient approach in terms of time cost.
[…]
3.3. RQ2: Characterizing Floating-Point Inconsistencies
@@ 3.3.1. Experimental Setup
Since LLM4FP outperforms the other LLM-based variants, we focus the rest of the evaluation on comparing LLM4FP with Varity with respect to the floating-point inconsistencies triggered by their generated floating-point programs. In this RQ, we discuss how often different classes of floating-point inconsistencies are triggered. We follow Varity to investigate five classes of floating-point inconsistencies: Real (normal and subnormal numbers), Zero (positive zero and negative zero), +Inf (positive infinity), -Inf (negative infinity), and NaN (not-a-number). Specifically, we count the number of times that there is a floating-point inconsistency of the kind $\{\textit{$r_{i}$},\textit{$r_{j}$}\}$, where $_{i}\neq r$j and both results belong to different numerical categories in $\{\textit{Real},\textit{Zero},\textit{+Inf},\textit{-Inf},\textit{NaN}\}$. For example, an inconsistency between a real number and a zero is counted once as a {Real, Zero } inconsistency.
@@ 3.3.2. Evaluation Results
Level Real, Real Real, Zero Real, +Inf Real, -Inf +Inf, -Inf
O0_nofma 720
O0 840 4
O1 830 4
O2 814 4
O3 798 4
O3_fastmath 1198 4 18 40 2
Total 5,280
@@
When further breaking down each inconsistency kind by optimization level for LLM4FP, we find that the $\{\textit{Real},\textit{Real}\}$ inconsistency class is observed across all optimization levels, as shown in Table 3. Among them, O3_fastmath contributes the largest number of inconsistencies. The extreme values from $\{\textit{Real},\textit{+Inf}\}$, $\{\textit{Real},\textit{-Inf}\}$, and $\{\textit{-Inf},\textit{+Inf}\}$ are relatively rare and mainly occur under O3_fastmath. The results show that the number of inconsistencies across optimization levels is relatively stable, except for O3_fastmath, which consistently amplifies inconsistencies.
[…]
Response to RQ2: The most common floating-point inconsistency triggered by the programs generated by LLM4FP is the $\{\textit{Real},\textit{Real}\}$ kind, which accounts for 98.48% of all inconsistencies and is observed across all optimization levels. In contrast, extreme-value categories involving +Inf, -Inf, or NaN are rare and occur in the most aggressive optimization level, O3_fastmath.
[…]
@@ 3.4. RQ3: Inconsistencies Across Compilers
In this RQ, we analyze whether numerical inconsistencies are more likely to occur between different host compilers or between host and device compilers, and how optimization levels influence these inconsistencies. We also analyze all possible combinations of compilers to see which combinations produce the largest and smallest differences. We consider the 16 hexadecimal digits of the floating-point result, and report an inconsistency if any of them do not match. Table 4 summarizes the inconsistency rates for LLM4FP and Varity across compiler pairs and optimization levels. In addition to the rates, it also reports the minimum, maximum, and average number of differing digits observed in inconsistent results.
@@
For both approaches, the highest inconsistency rates consistently happen in comparisons between host and device compilers, i.e., gcc vs. nvcc and clang vs. nvcc, rather than between host compilers such as gcc vs. clang. For example, in LLM4FP, the most inconsistent pair is gcc vs. nvcc, with a total inconsistency rate of 13.01%, compared to 3.74% for gcc vs. clang. A similar trend is observed in Varity, where the highest rate of 5.50% is from clang vs. nvcc, much higher than that of gcc vs. clang which is 1.17%.
@@
Optimization level plays an important role in when and how frequent numerical inconsistencies occur. While similar trends were observed in  Section 3.3, here we focus specifically on inconsistencies between host and device compilers. For Varity, the highest inconsistency rates consistently occur at the O3_fastmath level, reaching up to 2.06% for clang vs. nvcc, while the rates at other optimization levels remain very low, mostly below 1%. This is expected as aggressive optimizations and fast-math transformations can increase the chance of observing inconsistencies across compilers. For LLM4FP, the O3_fastmath level still yields the highest inconsistency rates across compiler pairs, peaking at 2.82% for clang vs. nvcc. However, unlike Varity, other optimization levels also exhibit relatively high rates around 2%, suggesting that LLM4FP triggers inconsistencies more broadly across compiler settings. This observation highlights the potential of LLM4FP for generating more complex and diverse floating-point programs that expose subtle compiler differences beyond aggressive optimization scenarios.
Beyond inconsistency rates, the digit difference results provide additional insight into the severity of numerical discrepancies. On average, LLM4FP inconsistencies exhibit smaller digit differences, i.e., from around 1 to 3 digits, compared to Varity, whose average differences are typically between 4 and 8 digits. This suggests that while LLM4FP exposes inconsistencies more frequently, most of them correspond to subtle numerical divergences rather than large-magnitude errors. In particular, LLM4FP mainly triggers inconsistencies of the $\{\textit{Real},\textit{Real}\}$ class, reflecting sensitive variations in normal floating-point computations rather than catastrophic precision loss.
[…]
@@ 3.5. RQ4: Comparing to O0 without FMA
Table 5. Inconsistency rates between any optimization level and O0_nofma for Varity and LLM4FP (higher inconsistency rates are preferred, and “–” indicates there is no inconsistency)
@@
As shown in Table 5, the inconsistency rate varies when comparing O0_nofma with different optimization levels. For Varity, the inconsistency rates at lower optimization levels from O0 to O3 remain near zero, all below 0.05%. It mainly detects differences caused by the most aggressive optimization level, O3_fastmath, especially for clang which shows an inconsistency rate of 2.07% compared to O0_nofma. In contrast, LLM4FP reports higher rates across all levels, with noticeable differences at moderate optimizations, reaching up to 0.78% for clang. The results indicate that LLM4FP’s programs are better at finding variations even when optimizations are less aggressive. For both Varity and LLM4FP, clang has the highest number of floating-point inconsistencies across its optimization levels, while nvcc is the most stable. Compared to Varity, LLM4FP finds more variations across optimization levels within each compiler.
[…]
Response to RQ4: Different optimizations within the same compiler can introduce different floating-point inconsistencies compared to -O0_nofma. LLM4FP finds inconsistencies across all optimization levels, not just the most aggressive ones. clang has the most inconsistencies while nvcc is the most stable.
[…]
@@ 4. Limitations
While LLM4FP improves the state of the art in generating floating-point programs that trigger numerical inconsistencies, it also has several limitations. First, our approach has a higher runtime cost due to the latency of LLM API calls which may limit scalability for large-scale testing. Nevertheless, the improved effectiveness of LLM4FP helps offset this overhead. Second, while we explicitly instruct the LLM to avoid undefined behavior, we cannot guarantee its absence, which may cause false positives in inconsistency detection. In the future, we plan to explore constraining our program generation rules and integrating runtime undefined behavior checkers such as GCC/Clang Undefined Behavior Sanitizer or AddressSanitizer (DBLP:conf/usenix/SerebryanyBPV12) to reduce false positives while maintaining valid program semantics.
Another limitation concerns prompt design and the interpretation of detected floating-point inconsistencies. The effectiveness of LLM4FP depends heavily on carefully crafted prompts. In practice, we find that good prompts combine structural and semantic constraints with explicit formatting rules and targeted mutation strategies, often supported by a successful example, whereas poor prompts tend to be vague or unconstrained, which can result in trivial or uninformative programs. If the prompts are poorly designed, the effectiveness of triggering numerical inconsistencies will be limited. Lastly, not all inconsistencies detected by LLM4FP indicate real bugs. Some may reflect allowed differences in compiler behavior, while others can impact program stability and lead to incorrect results. In either case, prior work (DBLP:conf/ipps/Laguna20; DBLP:conf/iiswc/SawayaBBGA17; DBLP:conf/sc/0007LR20; DBLP:conf/supercomputer/MiaoLR23; DBLP:conf/ics/MiaoLR24) has shown that it is important to be aware of floating-point inconsistencies across compiler and/or compiler optimizations.
@@ 5. Related Work
Compiler testing has been widely studied to ensure the correctness and reliability of compiler behavior (DBLP:journals/csur/ChenPPXZHZ20). Traditional compiler testing techniques include random testing (DBLP:journals/tse/DuranN84; DBLP:journals/cj/InceH86), randomized differential testing (DBLP:journals/dtj/McKeeman98; DBLP:conf/icse/GroceHJ07), equivalence modulo inputs testing (DBLP:conf/pldi/LeAS14), program enumeration (DBLP:conf/pldi/ZhangSS17) among others. The main challenge in compiler testing is to generate diverse, syntactically and semantically valid programs to test the compiler behaviors. Grammar-based approaches synthesize programs from scratch by following generation rules to ensure syntactic validity, such as CSmith (DBLP:conf/pldi/YangCER11), YARPGen (DBLP:journals/pacmpl/LivinskiiBR20), and LangFuzz (DBLP:conf/uss/HollerHZ12). To generate more realistic and complicated test programs, mutation-based approaches (DBLP:journals/pacmpl/DonaldsonELT17; DBLP:conf/pldi/LeAS14; DBLP:conf/oopsla/LeSS15; DBLP:conf/pldi/ZhangSS17) generate new programs by mutating valid seed programs.
@@
For floating-point programs, FLiT (DBLP:conf/iiswc/SawayaBBGA17) provides a test infrastructure for detecting variability in floating-point code across compilers and platforms. Unlike program generation approaches, FLiT does not synthesize new tests but instead systematically runs existing workloads to uncover cross-platform result differences. Varity (DBLP:conf/ipps/Laguna20) extends compiler testing to this domain by using grammar-guided random generation to synthesize test programs, focusing on numerical inconsistencies across different compilers and architectures rather than compiler bugs. However, it does not leverage domain knowledge or feedback from previous tests, which limits its ability to discover floating-point inconsistencies efficiently. In contrast, our work uses LLMs to generate and adapt test programs based on prior results, enabling us to discover more floating-point inconsistencies across compilers.
Several prior tools address different aspects of compiler-induced numerical inconsistencies but are complementary to our program-generation focus. pLiner (DBLP:conf/sc/0007LR20) isolates known numerical inconsistencies in C/C++ CPU programs at the line level using hierarchical code isolation and precision enhancement, while Ciel (DBLP:conf/supercomputer/MiaoLR23) extends it to heterogeneous programs and isolates inconsistencies at the expression level, enabling finer-grained analysis. Tools such as CIV (DBLP:conf/wcre/YuYYLCH23), S3FP (DBLP:conf/ppopp/ChiangGRS14), and CIGEN (DBLP:conf/ics/MiaoLR24) are input-generation methods that aim to expose floating-point inconsistencies or high-precision errors in existing programs, rather than synthesizing new programs. FPDiff (DBLP:conf/issta/VanoverDR20) performs differential testing on automatically identified synonymous functions across numerical libraries to detect inconsistencies in their results under specific inputs. Together, these approaches focus on isolating or exposing inconsistencies in existing programs or libraries, whereas our work targets automatic generation of new floating-point test programs to discover inconsistencies across compilers from scratch.
@@
Recent advances in LLMs have opened new opportunities for test program generation (DBLP:journals/corr/abs-2406-00515). LLMs like Codex (DBLP:journals/corr/abs-2107-03374) and GPT (DBLP:journals/corr/abs-2303-08774) have demonstrated strong capabilities in generating syntactically correct and semantically meaningful code from natural language prompts in tools like TestPilot (DBLP:journals/corr/abs-2302-06527) and ChatTester (DBLP:journals/corr/abs-2305-04207). Several recent works have applied LLMs to testing compilers (DBLP:conf/icse/XiaPTP024; DBLP:journals/pacmpl/YangDLY0J024; DBLP:journals/fgcs/MunleyJC24; DBLP:conf/sigsoft/Gu23). For example, WhiteFox (DBLP:journals/pacmpl/YangDLY0J024) integrates LLMs with compiler source code to guide white-box fuzzing for optimization testing. Fuzz4All (DBLP:conf/icse/XiaPTP024) uses LLMs to fuzz compilers across multiple languages via auto-prompting and iterative mutation. Unlike our work, they focus on general bug detection and do not target floating-point inconsistencies. Overall, our work bridges the gap between LLM-based program generation and numerical inconsistency detection across compilers, which is underexplored in prior research.